the value of tan(sin^-1 3/5+cos^-1 3/13 is equal to

Grade:12

1 Answers

Sher Mohammad IIT Delhi
9 years ago
5^2=3^2+4^2
13^2=12^2+5^2, the problem shouldtan(sin^-1 3/5+cos^-1 5/13), which can be rewritten as
tan(tan^-1 3/4+cos^-1 12/5); now use formula tan(a+b)=tan(a)+tan(b)/(1-tan(a)tan(b)), which gives
(3/5+5/13)/(1-3/5+5/13)
sher mohammad
b.tech, iit delhi